# Beyond Static Evaluation: A Dynamic Approach to Assessing AI Assistants’ API Invocation Capabilities

### v1

@@ Abstract
With the rise of Large Language Models (LLMs), AI assistants’ ability to utilize tools, especially through API calls, has advanced notably. This progress has necessitated more accurate evaluation methods. Many existing studies adopt static evaluation, where they assess AI assistants’ API call based on pre-defined dialogue histories. However, such evaluation method can be misleading, as an AI assistant might fail in generating API calls from preceding human interaction in real cases. Instead of the resource-intensive method of direct human-machine interactions, we propose Automated Dynamic Evaluation (AutoDE) to assess an assistant’s API call capability without human involvement. In our framework, we endeavor to closely mirror genuine human conversation patterns in human-machine interactions, using a LLM-based user agent, equipped with a user script to ensure human alignment. Experimental results highlight that AutoDE uncovers errors overlooked by static evaluations, aligning more closely with human assessment. Testing four AI assistants using our crafted benchmark, our method mirrored human evaluation with an correlation of $0.99$, marking an $8\%$ enhancement compared to conventional static evaluations.
@@
Honglin ${}^{\dagger}$, Yang ${}^{\dagger}$, Yunlong ${}^{\dagger}$
Xiaofeng ${}^{\ddagger}$, Yitong ${}^{\ddagger}$, Yutai ${}^{\ddagger}$, Wanxiang ${}^{\dagger}$thanks: ✉ Corresponding Author.
${}^{\dagger}$arbin Institute of Technology, ${}^{\ddagger}$ Huawei Technologies Co., Ltd.
@@ 1.   Introduction
Figure 2: An illustrative example for static evaluation, human evaluation and AutoDE, respectively. Sub-figure a shows the AI assistant correctly invoking an API call from a pre-defined dialogue history. In sub-figure b, the same assistant misses the “appName” parameter during human interaction, resulting in an incorrect API call. Sub-figure c demonstrates similar parameter issues when the assistant interacts with the user agent.
@@
By embedding task-specific information into the user script, the agent could reliably emulate human users. Our evaluations using different user agents not only showed a high correlation with human evaluations but also highlighted discrepancies when compared to static dialogue history evaluations. Specifically, our method leveraging Llama 2 7B Chat as the user agent attained $0.99$ consistency with human annotators in evaluating four AI assistants, among which Claude Instant 1.2 demonstrating the strongest performance. These findings underscore the need for, and the validity of, our dynamic evaluation approach.
@@ 3.   Method
Eval Method ICC3 R
GPT 3.5 $0.9866$ $0.9928$
Llama 2 7B Chat $\boldsymbol{0.9955}$ $\boldsymbol{0.9967}$
Static 0.9152915291529152 $0.9153$
Table 1: Comparison of consistency with human evaluations across various evaluation methods. “GPT 3.5” and “Llama 2 7B Chat” serve as user agents in AutoDE, whereas “static” refers to the static evaluation method. ICC3 and R refer to Intraclass Correlation Coefficient and Pearson Correlation Coefficient, respectively.
Assistant GPT 3.5 Llama 2 7B Chat Static Human
P R F1 P R F1 P R F1 P R F1
@@ 3.2.2.   User Script Generation
The dialogue context $S$ provides some level of background for the conversation, guiding the interactions between the user $\mathcal{U}$ and the assistant $\mathcal{A}$. Each user script corresponds to one usage scenario for the API. We directed GPT 4 OpenAI (2023) to brain-storm $5$ user script entries according to each API document to ensure diversity of the evaluation, resulting in a total of $330$ profiles. Each user script contains a dialogue context $S$ and an initial API call label $C_{g}^{\prime}$ which is later modified to a final version in section 3.2.3. A simplified example of such a dialogue context $S$ and API call label $C_{g}^{\prime}$ can be found in section 3.1.1 and section 2, respectively.
@@ 5.   Experimental Results
Going into more detail, Table 1 presents correlation metrics, specifically the Pearson Correlation Coefficient ($ICC3$) and the Pearson Correlation Coefficient $R$, for the different evaluation strategies. It’s noteworthy that the evaluation using Llama 2 7B Chat have a Pearson $R$ value that is $8\%$ higher than the static evaluation. This underlines that AutoDE offers a level of consistency with the human evaluation surpassing that of the static evaluation.

### v2

@@ Abstract
With the rise of Large Language Models (LLMs), AI assistants’ ability to utilize tools, especially through API calls, has advanced notably. This progress has necessitated more accurate evaluation methods. Many existing studies adopt static evaluation, where they assess AI assistants’ API call based on pre-defined dialogue histories. However, such evaluation method can be misleading, as an AI assistant might fail in generating API calls from preceding human interaction in real cases. Instead of the resource-intensive method of direct human-machine interactions, we propose Automated Dynamic Evaluation (AutoDE) to assess an assistant’s API call capability without human involvement. In our framework, we endeavor to closely mirror genuine human conversation patterns in human-machine interactions, using a LLM-based user agent, equipped with a user script to ensure human alignment. Experimental results highlight that AutoDE uncovers errors overlooked by static evaluations, aligning more closely with human assessment. Testing four AI assistants using our crafted benchmark, our method further mirrored human evaluation compared to conventional static evaluations.
@@
Honglin Mu, Yang Xu, Yunlong Feng
Xiaofeng Han, Yitong Li, Yutai Hou, Wanxiang Chethanks: ✉ Corresponding Author.
Harbin Institute of Technology, China
Huawei Technologies Co., Ltd., China
@@ 1.   Introduction
Figure 2: An illustrative example for static evaluation, human evaluation and AutoDE, respectively. Sub-figure A shows the AI assistant correctly invoking an API call from a pre-defined dialogue history. In sub-figure B, the same assistant misses the “appName” parameter during human interaction, resulting in an incorrect API call. Sub-figure C demonstrates similar parameter issues when the assistant interacts with the user agent. We demonstrate that certain API call issues related to interaction, concealed by static evaluation, can be revealed by dynamic human evaluation and AutoDE.
@@
By embedding task-specific information into the user script, the agent could reliably emulate human users. Our evaluations using different user agents not only showed a high correlation with human evaluations but also highlighted discrepancies when compared to static dialogue history evaluations. Specifically, our method leveraging Llama 2 7B Chat as the user agent attained higher consistency with human annotators in evaluating four AI assistants, among which Claude Instant 1.2 demonstrating the strongest performance. These findings underscore the need for, and the validity of, our dynamic evaluation approach.
@@ 3.1.1.   Manual Evaluation
Human Annotators
The annotators, two individuals with academic backgrounds in computer science, were trained to ensure a comprehensive understanding and execution of human-machine dialogue evaluation. Each system was annotated by one individual and reviewed for format and guideline consistency by a second. Prior to annotation, all personnel were required to read and understand a manual detailing the annotation tasks and procedures for handling anomalies, such as terminating dialogues in instances of multiple ineffective system responses.
@@ 3.2.2.   User Script Generation
The dialogue context $S$ provides some level of background for the conversation, guiding the interactions between the user $\mathcal{U}$ and the assistant $\mathcal{A}$. Each user script corresponds to one usage scenario for the API. We directed GPT 4 OpenAI (2023) to brain-storm $5$ user script entries according to each API document to ensure diversity of the evaluation, resulting in a total of $330$ profiles. Each user script contains a dialogue context $S$ and an initial API call label $C_{g}^{\prime}$ which is later modified to a final version in section 3.2.3. The prompt used to generate user script is shown as:
[system prompt]
You are an experienced prompt engineer.
[first round prompt]
Please construct 5 different use case
scenarios based on the following API
documentation:
{{API_DOC}}
Please follow the following format:
1.
Character: Lisa, a busy mother
Background: Lisa needs to take ...
Purpose: Using a tablet, Lisa books ...
API Call: {
    "funcName": "RegMedAppt",
    "time": "Monday",
    "departmentName": "Orthopedic"
InitialQuery: I want to book an medical
appoiment for next Monday at 1:30PM.
2.
Note that the generated scenarios  have
exactly five attributes...
where Character, Background, Purpose and API Call makes user script defined in (section 3.1.1). InitialQuery results in the initial query $H_{1}^{*}$ defined in section 3.1.1. The same set of examples is used across the construction.
@@ 3.2.3.   Static Dialogue History Generation
We addressed errors related to bad API understanding during static dialogue history generation. To mitigate them, we repeatedly invoked the model for correction, filtering out persistent inaccuracies to ensure test case accuracy.
@@ 4.2.   Assistant Model
Zephyr 7B Alpha
Zephyr is a series of language models that are trained to act as helpful assistants. Zephyr 7B Alpha is the first model in the series, and is a fine-tuned version of Mistral-7B-v0.1 Jiang et al. (2023). The model was fine-tuned on a variant of the UltraChat (Ding et al., 2023) dataset and further aligned with Direct Preference Optimization (DPO) (Rafailov et al., 2023) on the UltraFeedback (Cui et al., 2023) dataset.
@@ 5.   Experimental Results
Assistant GPT 3.5 Llama 2 7B Chat Static Human
P R F1 P R F1 P R F1 P R F1
Zephyr $48.08$ $50.16$ $48.69_{\pm 1.68}$ $49.76$ $51.21$ $50.05_{\pm 2.39}$ $80.69$ $79.77$ $80.01_{\pm 1.92}$ $48.70$ $50.26$ $49.14$
@@
Eval Method ICC3 R
GPT 3.5 $0.9869$ $0.9923$
Llama 2 7B Chat $\boldsymbol{0.9923}$ $\boldsymbol{0.9930}$
Static $0.8813$ $0.8813$
Table 2: Comparison of consistency with human evaluations across various evaluation methods. “GPT 3.5” and “Llama 2 7B Chat” serve as user agents in AutoDE, whereas “static” refers to the static evaluation method. ICC3 and R refer to Intraclass Correlation Coefficient and Pearson Correlation Coefficient, respectively. Note that this table is calculated with five data points and serves only as a reference.
Going into more detail, Table 2 presents correlation metrics, specifically the Pearson Correlation Coefficient ($ICC3$) and the Pearson Correlation Coefficient $R$, for the different evaluation strategies. It’s noteworthy that the evaluation using Llama 2 7B Chat have a Pearson $R$ value that is $11\%$ higher than the static evaluation. This underlines that AutoDE offers a level of consistency with the human evaluation surpassing that of the static evaluation.
@@ 8.   Ethics Statement
9.   Limitations
While promising, AutoDE has limitations to address. The benchmark dataset’s small scale warrants expansion for generalizability. Incorporating qualitative metrics beyond accuracy, like coherence and naturalness, could provide a more holistic evaluation. Evaluating comprehension of API documentation itself could reveal insights into reasoning abilities. Lastly, results on closed-source commercial models hinders reproducibility.
@@ 11.   Bibliographical References
  • Jiang et al. (2023) Albert Q. Jiang, Alexandre Sablayrolles, Arthur Mensch, Chris Bamford, Devendra Singh Chaplot, Diego de las Casas, Florian Bressand, Gianna Lengyel, Guillaume Lample, Lucile Saulnier, Lélio Renard Lavaud, Marie-Anne Lachaux, Pierre Stock, Teven Le Scao, Thibaut Lavril, Thomas Wang, Timothée Lacroix, and William El Sayed. 2023. Mistral 7b.
@@
  • Rafailov et al. (2023) Rafael Rafailov, Archit Sharma, Eric Mitchell, Stefano Ermon, Christopher D. Manning, and Chelsea Finn. 2023. Direct preference optimization: Your language model is secretly a reward model.
@@ 12.   Language Resource References
  • Cui et al. (2023) Ganqu Cui, Lifan Yuan, Ning Ding, Guanming Yao, Wei Zhu, Yuan Ni, Guotong Xie, Zhiyuan Liu, and Maosong Sun. 2023. Ultrafeedback: Boosting language models with high-quality feedback.